OpenStudy (anonymous):

find (dx/dy): (2x+3y)^5=x+1

OpenStudy (marco26):

I suggest using implicit differentiation to get dy/dx then get the reciprocal to find dx/dy.

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

why not just implicit x(y)?

OpenStudy (amistre64):

(2x+3y)^5=x+1 5(2x+3y)^4 (2x'+3) = x' and solve for x'

OpenStudy (anonymous):

I do not see a possible way to solve for x'

OpenStudy (amistre64):

let that 5()^4 part equal say: k k(2x'+3) = x' 2kx' + 3k = x' 3k = x' - 2kx' 3k = x'(1-2k) .... and divide

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

good luck :)